# Design-Space Exploration of SNN Models using Application-Specific Multi-Core Architectures

### 07 Feb 2024 v1

@@ 2. Project Description
Our project aims to analyze resource-efficient implementations of biologically inspired spiking neural networks, which on the one hand, enable the execution of SNNs in a resource-efficient manner and on the other hand, enable the possibility of online learning adaptation. The primary focus of the project is to explore the design space for potential computer vision applications (i.e., object detection/recognition). Artificial intelligence activities will employ a variety of applications, including one that uses an online learning strategy for object detection and recognition. In the first phase, we’$l$ study hardware architectures and a resource-efficient simulator that lets users run the model analyzing and visualization of the behavior using any reconfigurable computing platform such as CPU/GPU. Furthermore, our primary goal is to maximize performance by utilizing a CPU-based hardware platform using a multi-core architecture with configurable processors. Second, several neurons and synapse models will be assessed, along with their configurations. To perform in-silico analysis and simulation, it is extremely important to understand the different parameters of input-output concentration to obtain precise and accurate parametric values, i.e., what values are required to obtain a stable output concentration and which parametric values are dependent on one another in each event. Even with the moderate size of the set of neurons, defining and simulating the model using the combinations of these input concentrations may need highly precious balanced values. Alternatively, a run-time simulator offers a more useful method for simulating SNNs, allowing users to directly alter the simulation at any time to see how the model responds to changes in output concentration. This type of simulator not only provides an environment where users can completely comprehend this complicated structural model in an approachable way, but it also enables the SNN community to evaluate and visualize the model easily by raising or lowering the input at any level or instant in time. Also, it makes possible different learning processes that would not otherwise be easily possible.

### 25 Mar 2024 v2

@@ 2. Project Description
Our project aims to analyze resource-efficient implementations of biologically inspired spiking neural networks, which on the one hand, enable the execution of SNNs in a resource-efficient manner and on the other hand, enable the possibility of online learning adaptation. The primary focus of the project is to explore the design space for potential computer vision applications (i.e., object detection/recognition). Artificial intelligence activities will employ a variety of applications, including one that uses an online learning strategy for object detection and recognition. In the first phase, we’ll study hardware architectures and a resource-efficient simulator that lets users run the model analyzing and visualization of the behavior using any reconfigurable computing platform such as CPU/GPU. Furthermore, our primary goal is to maximize performance by utilizing a CPU-based hardware platform using a multi-core architecture with configurable processors. Second, several neurons and synapse models will be assessed, along with their configurations. To perform in-silico analysis and simulation, it is extremely important to understand the different parameters of input-output concentration to obtain precise and accurate parametric values, i.e., what values are required to obtain a stable output concentration and which parametric values are dependent on one another in each event. Even with the moderate size of the set of neurons, defining and simulating the model using the combinations of these input concentrations may need highly precious balanced values. Alternatively, a run-time simulator offers a more useful method for simulating SNNs, allowing users to directly alter the simulation at any time to see how the model responds to changes in output concentration. This type of simulator not only provides an environment where users can completely comprehend this complicated structural model in an approachable way, but it also enables the SNN community to evaluate and visualize the model easily by raising or lowering the input at any level or instant in time. Also, it makes possible different learning processes that would not otherwise be easily possible.